The perimeter of a triangle is 60cm. Find the length of the sides of the triangle which gives a maximum area.


1
Expert's answer
2022-01-30T16:20:33-0500

Let x, y, z be the sides of a triangle.

Then, the perimeter of the triangle is;

"\displaystyle
p(x,y,z)=x+y+z"

Also, using Heron's formula, the area of the triangle is;

"\displaystyle
A(x,y,z)=\frac{\sqrt{2x^2y^2+2x^2z^2+2y^2z^2-x^4-y^4-z^4}}{4}"


Now, from the question, "\displaystyle
P(x,y,z)=60". Thus, we are to;


maximize: "\displaystyle
A(x,y,z)=\frac{\sqrt{2x^2y^2+2x^2z^2+2y^2z^2-x^4-y^4-z^4}}{4}"

subject to: "\displaystyle
x+y+z-60=0"


Using Lagrange's method;


"\displaystyle
\Rightarrow\nabla\left(\frac{\sqrt{2x^2y^2+2x^2z^2+2y^2z^2-x^4-y^4-z^4}}{4}\right)=\lambda\nabla(x+y+z-60)\\
\Rightarrow\quad\frac{4xy^2+4xz^2-4x^3}{8\sqrt{2x^2y^2+2x^2z^2+2y^2z^2-x^4-y^4-z^4}}=\lambda\\
\qquad \frac{4yx^2+4yz^2-4y^3}{8\sqrt{2x^2y^2+2x^2z^2+2y^2z^2-x^4-y^4-z^4}}=\lambda\\
\qquad \frac{4zx^2+4zy^2-4z^3}{8\sqrt{2x^2y^2+2x^2z^2+2y^2z^2-x^4-y^4-z^4}}=\lambda\\
\quad\\
\text{Thus, we need to solve the following equations:}\\
\quad\\
4xy^2+4xz^2-4x^3=\lambda8\sqrt{2x^2y^2+2x^2z^2+2y^2z^2-x^4-y^4-z^4}\ \cdots\cdots\cdots\cdots\cdots\cdots(i)\\
4yx^2+4yz^2-4y^3=\lambda8\sqrt{2x^2y^2+2x^2z^2+2y^2z^2-x^4-y^4-z^4}\ \cdots\cdots\cdots\cdots\cdots\cdots(ii)\\
4zx^2+4zy^2-4z^3=\lambda8\sqrt{2x^2y^2+2x^2z^2+2y^2z^2-x^4-y^4-z^4}\ \cdots\cdots\cdots\cdots\cdots\cdots(iii)\\
x+y+z-60=0\cdots\cdots\cdots\cdots\cdots\cdots(iv)"


substituting (i) into (ii) and (iii) yields;


"\displaystyle
4yx^2+4yz^2-4y^3=4xy^2+4xz^2-4x^3\\
4zx^2+4zy^2-4z^3=4xy^2+4xz^2-4x^3\\
x+y+z-60=0\\
\Rightarrow\\
yx^2+yz^2-y^3-xy^2-xz^2+x^3=0\\
zx^2+zy^2-z^3-xy^2-xz^2+x^3=0\\
x+y+z-60=0\\
\Rightarrow\\
xy(x-y)-z^2(x-y)+(x^3-y^3)=0\\
xz(x-z)-y^2(x-z)+(x^3-z^3)=0\\
x+y+z-60=0\\
\quad\\
\text{But }(a^3-b^3)=(a-b)(a^2+ab+b^2)\\
\Rightarrow\\
xy(x-y)-z^2(x-y)+(x-y)(x^2+xy+y^2)=0\\
xz(x-z)-y^2(x-z)+(x-z)(x^2+xz+z^2)=0\\
x+y+z-60=0\\
\Rightarrow\\
(x-y)[xy-z^2+(x^2+xy+y^2)]=0\\
(x-z)[xz-y^2+(x^2+xz+z^2)]=0\\
x+y+z-60=0\\
\Rightarrow\\
(x-y)[x^2+x(2y)+(y^2-z^2)]=0\\
(x-z)[x^2+x(2z)+(z^2-y^2)]=0\\
x+y+z-60=0\\
\Rightarrow\\
(x-y)(x+y-z)(x+y+z)=0\\
(x-z)(x+z-y)(x+z+y)=0\\
x+y+z-60=0\\
\Rightarrow\\
(y-z)\times z\times(2y+x)=0\\
2y+z-60=0,\ \text{if }x=y\\
\Rightarrow\\
3z-60=0,\ \text{if }y=z\\
\Rightarrow\\
x=y=z=20"


Hence, the length of the sides of a triangle of perimeter 60cm which gives a maximum area are;

"\displaystyle
x=y=z=20\ cm", where x, y, z are the sides of the triangle.